OpenStudy (anonymous):

(3x^2)3x^2 (3x^3)^2(x^2) Which one has the largest exponent when its simplified?

OpenStudy (turingtest):

use the rules\[(x^a)^b=x^{ab}\]and\[x^a\cdot x^b=x^{a+b}\]

OpenStudy (anonymous):

can u help me solve it?

OpenStudy (turingtest):

\[(3x^2)\cdot(3x^2)=(3\cdot3)(x^2\cdot x^2)\]now apply the rules above